UCL Ear Institute


16 January 2014 Minutes

Health and Safety Meeting held on 16th January 2014.


  • Dr S Dawson (SD)

  • Mr G Nevill (GN)

  • Dr N Daudet (ND)

  • Miriam Gomez

  • Mrs C Overington (CO)

  • Mr W Potter (WP)

  • Mr A Pyne (AP)

  • Ms J Ziff (JZ)


  • Dr P Singh (PS)


  • Actions from last meeting - GN to chase company who are fixing the oxygen sensors. GN confirmed completed (£600)


  • Black Gothic Church door locked during fire alarm test – AP confirmed work completed and passed new fire test.
  • Build-up of glass bottles in Lab: company who supply the bottles to be contacted re do they collect the bottles, MG confirmed completed. Recycling in general Estates are chasing up issue of glass collection (2nd floor lab and around the building). There will be general bins, the contents of which will be separated off site (see AOB).
  • Safety notices: AP confirmed completed putting up safety notices. GN confirmed notices have correct information.
  • Chemical Waste collection: Batteries are being taken to Tesco or Post office to be disposed of, currently collections building in BP’s room and MG’s room.
  • Service Contract for main autoclave – MG has renewed the contract. AP to find out whether UCL central are financially liable for the contract – AP confirmed it is central’s responsibility. MG confirmed the issue has been chased this week, AP suggested MG chase Gary Stratmann.
  • Keypad on Jonathan Ashmore’s door – AP to replace keypad. This has not been actioned as Jonathan stated it is not needed. SD concerned it is an issue as keypad entry in all safety and ethics documents. AP stated one entrance to that lab is without a keypad. SD concerned of the impact this would have on radioactive inspections. ND also highlighted issue with keypad on 1st floor which was confirmed by SJ. AP confirmed this is an outstanding issue from when EI was transferred from central estates responsibility to Satellite Estates responsibility.
  • First Aiders – Clarification required for the first aid list. GN asked Kate to produce updated list. GN confirmed lists are all up to date.


  • Shelving insufficiently stored – AP to include in inspection list and remove. AP confirmed removed.
  • Blue Gloves – BP confirmed blue gloves are still being worn in non-laboratory areas, not as much as they were. Discussion had regarding the hazard to others and also the hazard to person handling the materials as not always practical to remove the gloves.
  • Restrictions to office visibility on 1st floorAP confirmed he arranged black out blind, however functional rather than smart.


  • GN queried whether Soundseekers (SS) should be part of Safety Committee. All agreed that they need not, unless a specific issue was raised. However SS staff should be aware of the building health and safety procedures and have appropriate safety officers within their team.


  • New student rep required – completed as SJ has joined the committee.


  • AP to undertaken monthly inspections – SD and AP reiterated the need for monthly inspections. AP not completing monthly at present, see item Safety Office Update for further discussion.


  • MG to include Gary Strattman on emails relating to autoclave
  • MG and BP to combine battery collections and contact Estates to arrange regular collections
  • AP will put in request for new lock
  • GN to send reminder email for staff to remove blue gloves before entering non laboratory areas
  • AP to liaise with Lucy / Stuart within SS to suggest they contact GN to arrange for standard staff walk around safety induction. SS can then complete training for their team.



  • AUTOCLAVE - passed tests.
  • AIR CONDITIONING IN LABORATORY - over 25° at present. AP confirmed the fault with fan coincides with the changing of the filter.
  • HOOD NOT WORKING IN PATHOLOGY ROOM - GN confirmed reported twice. SD raised concern that there always seemed to be one fume hood not working. Discussion had regarding best way to report such issues. AP confirmed these issues should be reported to satellite estates helpdesk website and cc AP into the email. ND confirmed the helpdesk email is available under ‘staff resource’ on the EI website.
  • FLOOR -80 FREEZER ROOM - flooring is secured with only tape but requires full replacement.


  • AP send email to relevant staff regarding how to report issues i.e. via satellite estates email instead of online.
  • AP to look into faulty floor in freezer room and add to issues to be raised with Satellite Estates


  • STEVE TIDMARSH IS THE NEW UCL SAFETY SERVICES, AREA SAFETY OFFICER - GN has not met him yet but extended an open invitation to visit EI.
  • FIRE MARSHALL ISSUES – Katie Smith and Nick Hayward now trained. The human areas on 1st floor are still underrepresented. Top floor DV and IY only fire marshalls, use of room going back to PhD student shortly, now EviDENT moved to Hospital. Appropriate fire marshal depends on which team moves into the area, GN suggested Kim Airey be responsible for EviDENT fire marshal issues. Their fire rules relate to the hospital and comes under their lease. 3 members of EviDENT team (likely Kim, Helen and one other) are required to be fire marshalls. Skills lab still needs fire marshal representative. Jose is a fire marshal already. People in skills lab changes a lot so it is difficult to identify an appropriate person and students are in there unsupervised.
  • FIRE MARSHALL TRAINING – Training is due to alter, to be delivered on site for site specific training. Therefore any retraining due would be better to wait until on-site training scheduled for 4th September 2014. There has been a request for this date to be altered as it coincides with staff holiday period.
  • FULL SITE FIRE ASSESSMENT 3rd SEPTEMBER 2014 - Access to all areas required, therefore all keys and appropriate staff to be available. Also need discussion and relevant records i.e. PAT Testing, Fire Marshall training and fire risk assessment records. SD suggested showing Safety Committee Meeting minutes too.
  • CHAMBER EXIT - Door now locked, AP and BP confirmed no further action required.
  • PAT TESTING – Every 5 years. AP completed last time and now not due here for a while. EviDENT team will need their equipment PAT tested.
  • MONTHLY CHECKLIST – GN and AP discussed what tasks should be included. Discussion had with regards to whether the tasks should be shared out, concern that this might lead to the tasks not all being completed, important that AP knows for himself what has actually been completed.
  • SAFETY POLICY ACCESS – Some confusion as to whether the Safety Policy is accessible on the website. Minutes from last Safety Committee also not available on the site.
  • DEAF MESSAGE SERVICE – This has yet to be installed. Text comes through to deaf member of staff’s phone if the fire alarm goes off. Currently not linked to central panel. Also concern that lack of alarm sounder on the booths. AP has reported, Estates Manager is contacting FISK.


  • GN to ask Vit to become fire marshal to cover area not covered on 1st floor
  • AP to arrange PAT Testing for EviDENT equipment
  • AP and GN to adapt UCL standard monthly checklist for EI purposes and start undertaking the checking
  • ND emailed KF during meeting to request Safety Policy be available to all staff
  • AP to see whether the fire alarm was missed off when installed in booth in 2013


  • RISK NET – up and running re risk assessment forms. Two have gone through so far (Jennifer Linden and Jennifer Bizley’s).
  • CONSULTED J BLACKMAN ON RENEWAL OF RISK ASSESSMENTS – does it need renewing, maybe every 5 years? Revise risk assessment once Risknet up and running. Not necessarily a resubmission. ND contacting people who have active projects at present, double checking the list is up to date. All emails from GM go direct to ND, then forwards only when appropriate i.e. new regulations etc.


  • ND will contact again to get more information and clarification.


  • No usage for some time. Room is now kept locked as requested by environment agency on last visit.


  • EviDENT TEAM MOVE - AP confirmed EviDENT team have moved to the hospital. Team agreed that as EviDENT team are located in hospital they do not need a member of their staff to represent them at the Safety Committee.
  • LEAK – within finance office. AP met with satellite facility manager and asked to chase up.
  • CLEANING – improving in laboratories. SD has noticed cobwebs, AP explained there are issues re cleaners working at height. ONG have won the cleaning contract for entire UCL. CO confirmed there is a general lack of cleanliness including soap dispensers not being filled etc. AP suggested any staff shortage causes lack of section to be completed. General office cleanliness not great, but they are limited in what they can do. SJ confirmed there had been issues with the cleaner removing items from desks which they should not have done and also general rubbish not being collected i.e. food wrappers on stairs. Team discussed clean desk policy but agreed cleaners are still likely to not clean desks as the contract is UCL wide. AP has tried to clarify their routine / duties.
  • SEAL ON LIQUID NITROGEN TANK FAULTY – MG confirmed lid recently replaced. GN suggested replace whole seal. SD requested a replacement from grants. On Cathy’s list for people to put on grant as communal piece of equipment. Smaller self-automating option now available.


  • AP to request the Service Level Agreement for ONG.


  • CO and Prof Saeed thought it important to discuss the HTA at the Safety Committee meeting. Prof Saeed leads and completes the document, CO meeting with Prof Saeed re section 20 and other sections which relate to safety. Majority covered by UCL policy. It is completed annually. Prof Saeed is the licence holder. Useful for staff to know it exists and where to access the information, although this is restricted.


  • Safety Committee Members to ensure necessary staff aware of existence of HTA
  • CO to give GN copy of HTA
  • CO to request HTA is uploaded to website within secure area (not for general site)


  • All agreed that this was fully covered. PS has not been available to attend the last few meetings.


  • SD to ask PS to nominate a teaching colleague or Robert Heller to attend Safety Committee meetings on PS’s behalf


  • GREEN IMPACT ISSUES - MG confirmed the laboratories have yet to be registered for Green Impact Issues.
  • AP explained delay with recycling bins is due to the lids not being provided and the disagreement between departments as to who should pay for the lids. Once they arrive all other bins (other than clinical waste) will be removed. The rubbish will be sifted off site. Currently all cardboard should be left outside the kitchen to be recycled. Some general waste is going into yellow bags incorrectly at great cost.


  • MG to register lab areas with Green Impact Issue
  • AP to get clear bags as stop gap until the recycling lids arrive.

NEXT MEETING: To be held in approximately 6 months